263 dogs entered at a dog show have an average weight of 11.9 pounds with a standard deviation of 8.3 pounds. 7 dogs are selected at random. Assuming a normal distribution, what is the probability that their average weight is between 7.1 and 21.5 pounds?
Answer #1

solution:-

given that mean = 11.9 , standard deviation = 8.3

7 dogs are selected at random so n = 7

=> P(7.1 < x < 21.5) = P((7.1-11.9)/(8.3/sqrt(7)) < z < (21.5-11.9)/(8.3/sqrt(7)))

= P(-1.53 < z < 3.06)

= P(z < 3.06) - P(z < -1.53)

= P(z < 3.06) - (1-z < 1.53)

= 0.9989 - (1-0.9370)

= 0.9359